$SPAD/src/input parpcurv.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

\getchunk{license}
-- test for ParametricPlaneCurve domain
draw(curve(sin(3*t),cos(3*t)),t = -5*%pi..5*%pi)
draw(curve(sin(tan(t)),tan(sin(t))),t = -5*%pi..5*%pi)

———-

References^{[1] nothing}

4